LOW-FREQUENCY VIBRATIONAL MODES OF ALUMINUM-PHOSPHINE AND -ARSINE COMPLEXES

SHENGGANG LI, GRETCHEN K. ROTHSCHOPF AND \underline{DONG-SHENG YANG}, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Vibrationally resolved spectra of the aluminum-phosphine and -arsine complexes were measured using ZEKE spectroscopy. The ZEKE spectrum of Al-P(CH_{3})_{3} reveals the vibrational energy levels of low-frequency modes in the cationic and neutral species. These modes include the symmetric Al^{+}-P stretching (326 cm^{−1}) and Al^{+}-P-C bending (76 cm^{−1}) in the ion and the symmetric Al-P-C bending (71 cm^{−1}) in the neutral. The spectrum of Al-As(CH_{3})_{3} displays a major progression associated with the Al^{+}-As-C bending. Other transitions are also evident in the spectrum. Analysis is ongoing.